### v1

# Multi-Agent Reinforcement Learning for Connected and Automated Vehicles Control: Recent Advancements and Future Prospects

Abstract
Connected and automated vehicles (CAVs) have emerged as a potential solution to the future challenges of developing safe, efficient, and eco-friendly transportation systems. However, CAV control presents significant challenges, given the complexity of interconnectivity and coordination required among the vehicles. To address this, multi-agent reinforcement learning (MARL), with its notable advancements in addressing complex problems in autonomous driving, robotics, and human-vehicle interaction, has emerged as a promising tool for enhancing the capabilities of CAVs. However, there is a notable absence of current reviews on the state-of-the-art MARL algorithms in the context of CAVs. Therefore, this paper delivers a comprehensive review of the application of MARL techniques within the field of CAV control. The paper begins by introducing MARL, followed by a detailed explanation of its unique advantages in addressing complex mobility and traffic scenarios that involve multiple agents. It then presents a comprehensive survey of MARL applications on the extent of control dimensions for CAVs, covering critical and typical scenarios such as platooning control, lane-changing, and unsignalized intersections. In addition, the paper provides a comprehensive review of the prominent simulation platforms used to create reliable environments for training in MARL. Lastly, the paper examines the current challenges associated with deploying MARL within CAV control and outlines potential solutions that can effectively overcome these issues. Through this review, the study highlights the tremendous potential of MARL to enhance the performance and collaboration of CAV control in terms of safety, travel efficiency, and economy.
Index Terms:
Connected and automated vehicles, Multi-agent reinforcement learning, Intelligent transportation systems, Vehicle control.
I Introduction
The transportation industry is currently experiencing a significant revolution, driven by the utilization of advanced techniques such as artificial intelligence (AI), the Internet of Things (IoT), and sensor technologies [1].
Notably, automated vehicles (AVs) represent a pivotal development in this ongoing transportation revolution with their potential to alleviate traffic congestion, improve road safety, and enhance accessibility [2]. Recently, the advent of connected and automated vehicles (CAVs) has taken this revolution a step further, combining the capabilities of AVs to establish a networked transportation ecosystem by using advanced communication technologies, such as vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-cloud (V2C) [3, 4]. This interconnection and collaboration significantly augment the capabilities of individual AVs by extending their perception beyond immediate line-of-sight constraints and enhancing their collaborative decision-making abilities, thereby improving safety and traffic efficiency [5, 6, 7, 8]. However, the development of a robust CAV paradigm that tightly integrates computing, communication, and control technologies presents a considerable challenge, especially in the presence of complex interactions among AVs and intricate traffic scenarios involving other road users (e.g., bicyclists, pedestrians, and human-driven vehicles (HDVs)).
I-A Development of CAV control
Control of CAVs stands as a key process within the above revolution and enables data-driven enhancements to transportation safety and sustainability by integrating seamlessly with diverse disciplines such as energy management, urban planning, and the broader social context [9, 10, 11]. Conventional optimization control methods have been proposed to tackle the CAV control problems [12, 13]. Specifically, an idealized model predictive control (MPC) approach is introduced in [14] to coordinate the behaviors of all following CAVs in a platoon, resulting in minimal control delay and efficient management of car-following behaviors to mitigate traffic oscillations. In [13], the coordination and control of the CAVs at road intersections is formulated as a non-convex optimal control problem and solved through semidefinite relaxation with randomization. However, the effectiveness of these approaches relies heavily on accurate system modeling [15], which is often not available. Furthermore, these methods commonly require online optimization that demands substantial computational resources that render these approaches impractical for real-time applications of CAV control [16].
On the other hand, reinforcement learning (RL), a data-driven approach, has gained substantial attention within the research community due to its outstanding abilities in addressing sequential decision-making problems, such as gaming [17], robotics [18], behavioral planning [19, 20, 21, 22], intelligent energy management [23, 24]. Similarly, the AV control has begun exploring the potential of employing RL for various traffic scenarios [25, 26, 27]. For instance, in [26], a safe RL framework based on an enhanced double deep Q-Network (DDQN) [28] is proposed for highway lane-changing, resulting in zero collisions. In [25], an RL-based car-following model is introduced to optimize driving behavior, enhancing travel efficiency, fuel economy, and safety at signalized intersections in real-time performance. This approach featuring an effective reward function, demonstrates strong performance under varying traffic demands and traffic light cycles with different durations. Taking into account the role of HDVs in the control of AVs, Shi et al. employ a distributed proximal policy optimization (DPPO) control strategy that allows them to learn from and respond to the behavior of the HDVs, optimizing performance at both the local subsystem level and the broader mixed traffic context [29]. Qu et al. introduce a control approach utilizing the deep deterministic policy gradient (DDPG) algorithm aimed at reducing traffic fluctuations and enhancing fuel economy. However, their approaches only consider a single agent (i.e., AV) and exhibit poor generalization when multiple agents are involved. Additionally, their study focuses solely on individual vehicle control and does not leverage shared information [30].
I-B Multi-agent reinforcement learning for CAV control
In scenarios involving multiple agents, multi-agent reinforcement learning (MARL), a natural extension of single-agent RL, has emerged as a promising tool for optimizing the collective behaviors of multiple agents, showing significant potential for advancing the control strategies employed in the CAV ecosystem [31, 32, 33, 34, 35]. Fig. 1 illustrates a comprehensive multi-agent system, i.e., CAVs, and depicts a holistic approach to CAV operation, from sensory input and data processing to inter-agent communication and eventual decision-making and motion execution. On the left, the primary input sources include a variety of sensors such as GPS/INS, Camera, Radar, and Lidar. These sensors play a vital role in gathering real-time environmental data. The data is then processed through sensor fusion, which integrates multiple data sources to produce a consistent, accurate, and useful description of the environment. This processed data is crucial for tasks like localization (determining the vehicle’s precise position), perception (recognizing objects and obstacles), and environment prediction (forecasting possible changes in the surroundings). Furthermore, the system is equipped to use historical data and maps, ensuring that the vehicle’s decision-making is informed not only by real-time data but also by past experiences and pre-existing geographic information.
Figure 1: The development of connected and autonomous vehicles
The system’s architecture is anchored by multiple AVs, each endowed with its own repository for private data as well as access to a pool of shared information. These AVs are equipped with the capability to engage in dialogue and data exchange with neighboring units via local communication links, empowering them to enhance their decision-making processes and broaden their situational awareness. Next to this, the decision-making and motion planning module delineates the system’s logical and strategic framework. Based on the rich data inputs and dynamic interplay of inter-agent communications, decisions are formulated and motion trajectories are charted. The steering control and powertrain control modules demonstrate how these decisions translate into physical actions, such as turning the steering wheel or adjusting the engine’s or motor operation. A comprehensive examination and discourse on these components will be presented in Sections 3 and 4.
Take some instances for example, a fully scalable and decentralized MARL algorithm is developed in [32] for large-scale traffic signal control, demonstrating superior control performance over the other state-of-the-art decentralized MARL algorithms. Furthermore, in [35], a scalable and safe MARL algorithm based on the parameter-sharing scheme is proposed for on-ramp merging in mixed traffic. Despite recent progress, the development of efficient and scalable MARL algorithms still presents significant challenges. The primary concern is guaranteeing the robustness and safety of these algorithms, especially in the face of diverse and uncertain traffic scenarios [35]. Furthermore, the dynamic behaviors of CAVs lead to non-stationary environments, adding layers of complexity to the development of MARL algorithms [32]. Another challenge lies in the lack of realistic simulators capable of accurately modeling traffic scenarios and other road users, which are essential for effective training of MARL algorithms [34]. Lastly, the real-world implementations of these algorithms remain an uphill task due to sim-to-real gaps [36, 35], and technological, legislative, and societal barriers [37]. While a few articles describe the potential of MARL in CAVs [38, 39, 40], none of them, to the best of our knowledge, are specifically devoted to the application of MARL for CAV control in intelligent transport systems (ITS) domain.
I-C Contributions of this review
This paper is thus intended to provide a comprehensive and systematic review of MARL in the context of CAVs. We survey the recent developments in MARL algorithms, discussing their diverse applications in aspects of CAV control based on the extent of control dimensions, such as platooning control, lane changing, on-ramp merging, and traffic signal cooperation. We also aim to highlight and delve into the technical challenges that MARL faces in the context of these applications and discuss potential research directions that could be pursued to address these challenges. This review not only serves to provide a clear, up-to-date understanding of the current state-of-the-art MARL algorithms applied in CAV control, but it also outlines the potential direction for future work in this area. The main contributions of this paper are summarized as follows.
  1. 1.
    A self-contained overview of RL and MARL background, specially tailored for the automotive community, is provided, recognizing that these concepts may not be widely familiar within this specific domain.
  2. 2.
    We review recent advances in MARL for CAV control, proposing a new taxonomy of MARL algorithms based on the degree of control signals. Furthermore, we provide an extensive examination of the leading simulation platforms employed in MARL research for CAVs.
  3. 3.
    We discuss the remaining challenges and potential future directions, shedding light on the exciting opportunities that lie ahead in this field.
  4. 4.
    We establish a well-classified list of relevant papers, which will be actively updated and can be accessed at the following site: https://github.com/huahuaedi/MARL_in_CAV_control_review.
The rest of this paper is organized as follows: Sec. II provides the preliminaries of RL and MARL; a detailed review of MARL applied in CAV control from different control dimensions and mainstream simulation platforms is described in Sec. III. Sec. IV summarizes the remaining challenges and opportunities. Finally, Sec. V concludes the review.
II Backgrounds
In this section, we offer an in-depth overview of the core principles of RL. Subsequently, we explore several leading MARL algorithms, establishing the context to enhance the understanding of our review.
II-A Preliminaries of reinforcement learning (RL)
RL, often mathematically formulated as a Markov Decision Process (MDP), has emerged as a promising data-driven method for sequential decision-making [41]. Recent advancements in Deep Neural Networks (DNNs) have further bolstered RL’s capabilities in handling intricate tasks. Noteworthy examples of these algorithms include the deep Q-network (DQN) [42], deep deterministic policy gradient (DDPG) [43], and advantage actor-critic (A2C) [44]. For instance, the AlphaGo computer program, based on DQN, made history by becoming the first of its kind to defeat a professional human player in Go and subsequently beat a world champion in the game [42]. Additionally, in a study by Lee et al. [18], researchers successfully applied Trust Region Policy Optimization (TRPO, [45]) to navigate quadrupedal robots across challenging terrains, including surfaces like mud and snow, and through dynamic footholds.
In an RL framework (refer to Figure 2), the learner (referred to as the agent) traverses the environment employing a trial-and-error methodology. The agent makes decisions, executes corresponding actions within the environment, and, in turn, receives a reward signal along with a new state. This reward, conveyed by the environment, offers feedback to the agent, indicating whether the impact of its actions is positive or negative. Mathematically, the RL problem can be formulated as a Markov Decision Process (MDP). The MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ is defined as follows:
  1. 1.
    State space $\mathcal{S}$: A comprehensive set of states describing the environment, specifying the position and conditions of agents at a given time step $t$. An observation $o$ provides a partial state description and may lack complete information. The environment is fully observed if the agent can access the entire state; otherwise, it is partially observed.
  2. 2.
    Action space $\mathcal{A}$: A collection of valid actions within the environment. Environments, such as Atari and Go, have discrete action spaces, limiting the agent to a finite set of moves. Environments controlling a physical robot have continuous action spaces, where actions are real-valued vectors.
  3. 3.
    Reward $\mathcal{R}(s_{t},a_{t},s_{t+1})$:The environment returns a reward upon executing action $a_{t}$ at state $s_{t}$. The reward signal’s value may be positive or negative, depending on the agent’s actions.
  4. 4.
    Transition Probability $\mathcal{P}_{ss^{\prime}}(S_{t+1}=s^{\prime}|S_{t}=s)$: the transition probability describes the likelihood of an agent moving from one state to another.
Figure 2: Illustration of reinforcement learning (RL).
In the subsequent subsections, we will explore three prevalent RL algorithms: Deep Q-learning, Policy Gradient, and the Actor-Critic Network.
II-A1 Deep Q-Learning
In Q-learning, the Q-function, denoted as $Q_{\theta}$, is often parameterized by a set of parameters $\theta$. This parameterization can be achieved using various function approximators, such as Q-tables [46], Linear Regression (LR) [47], and Deep Neural Networks (DNNs) [42]. The temporal difference, denoted as $(\mathcal{T}Q_{\theta^{-}}-Q_{\theta})(s_{t},a_{t})$, forms the basis for updating $\theta$. Here $\mathcal{T}$ and $Q_{\theta^{-}}$ represent the dynamic programming (DP) operator and a recently frozen model $\theta^{-}$ [4], respectively. To reduce variance in estimating Q-values and improve exploration, techniques like the $\epsilon-\text{greedy}$ and experience replay are commonly integrated into deep Q-learning [47]. We can obtain the optimal action $a^{*}(s)$ through the $Q^{*}(s_{t}=s,a_{t}=a)$ as:
$a^{*}(s)=\operatorname*{arg\,max}_{a}{Q^{*}(s_{t}=s,a_{t}=a)},$ (1)
Some widely recognized deep Q-learning algorithms include DQN [48], double deep Q-Network (DDQN)[28], and hindsight experience replay (HER) [49].
II-A2 Policy Gradient
In contrast to the Q-learning approach, the policy gradient method directly parameterizes the policy, represented as $\pi_{\theta}$, using a set of parameters $\theta$. The objective of updating $\theta$ is to enhance both the likelihood of taken actions and the cumulative rewards. This is accomplished through the loss function:
$\nabla_{\theta}\mathcal{L}(\pi_{\theta})=\mathop{\mathbb{E}}_{\tau\sim\pi_{\theta}}[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R_{t}],$ (2)
Subsequently, the policy network parameters undergo an update through stochastic gradient ascent, expressed as:
$\theta_{k+1}=\theta_{k}+\alpha\nabla_{\theta}\mathcal{L}(\pi_{\theta}),$ (3)
Compared to Q-learning, policy gradient methods are robust to non-stationary transitions within each trajectory. However, it’s essential to note that they tend to exhibit high variance [32]. Renowned algorithms utilizing policy gradients include Trust Region Policy Optimization (TRPO) [45], Proximal Policy Optimization (PPO) [50], and Deep Deterministic Policy Gradient (DDPG) [43].
II-A3 Actor-critic Network
To address the high variance associated with sample returns in the policy gradient method, actor-critic algorithms, such as A2C [44], incorporate the advantage function to refine the policy gradient technique, leveraging both policy (actor) and value (critic) functions. The advantage function is represented as:
$A^{\pi}(s_{t},a_{t})=Q^{\pi_{\theta}}(s_{t},a_{t})-V_{w}(s_{t}),$ (4)
Here, the parameters $\theta$ are updated through the policy loss function:
$\nabla_{\theta}\mathcal{L}=\mathop{\mathbb{E}}_{\pi_{\theta}}[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})A_{t}],$ (5)
Simultaneously, the value function updates with:
$\mathcal{L}=\min_{w}\mathop{\mathbb{E}}_{\mathcal{D}}[(R_{t}+\gamma V_{w^{-}}(s_{t})-V_{w}(s_{t}))^{2}],$ (6)
where $\mathcal{D}$ represents the experience replay buffer, aggregating past experiences. This buffer works alongside parameters derived from previous iterations, typically used in a target network [35].
However, despite significant advancements, single-agent RL often struggles with scalability challenges, particularly in real-world control scenarios involving multiple agents. These challenges stem from inherent non-stationarities and the partial observability intrinsic to such systems [33].
II-B Multi-agent reinforcement learning (MARL)
In multi-agent systems involving multiple agents (e.g., AVs, see Fig. 3), the agent’s actions affect not only their own rewards but also the rewards of other agents. MARL has found great success across a wide range of multi-agent systems, including traffic light control[32], games [51], resource management in wireless networks [52], and power grid control [53], only to name a few. A prominent and simplistic methodology in this domain is Independent Q-learning (IQL) [54]. In IQL, each local Q-function predominantly relies on the local action, represented as $Q_{i}(s,a)\approx Q_{i}(s,a_{i})$. Along similar lines, the Independent Advantage Actor-Critic (IA2C) serves as an actor-critic version of MARL, as proposed by [4]. While both IQL and IA2C offer highly scalable solutions, they grapple with challenges posed by partial observability and non-stationary Markov Decision Processes (MDP). This is mainly attributed to their intrinsic assumption: the behaviors of all other agents are perceived as part of the environment’s dynamics. This becomes problematic considering these agents’ policies take continual updates during training [32].
[…]
Figure 3: Illustration of multi-agent reinforcement learning (MARL).
II-C Training and execution strategies in MARL
In MARL settings, different approaches and frameworks are employed to tackle the complex challenges of coordinating and training agents in decentralized environments. This subsection explores two fundamental paradigms: Centralized Training with Decentralized Execution (CTDE) and Decentralized Training with Decentralized Execution (DTDE), each offering distinct insights and trade-offs in the pursuit of effective multi-agent learning [56, 57, 58].
II-C1 Centralized training with decentralized execution (CTDE)
The majority of MARL algorithms adhere to centralized training with a decentralized execution (CTDE) framework. In this framework, the decentralized problem is initially transformed into a centralized one, solvable by a central controller, which gathers essential training information, e.g., observation, reward, and global state information, for the training process. Following this data collection, the centralized value functions are learned based on the information of all agents, and then the gradient from the centralized value function is used to train the policy of each agent. But in the process of execution, each agent outputs action according to its individual observation (see Fig. 4).
It is important to acknowledge the inherent trade-offs of the CTDE. One significant strength is the enhanced learning stability enabled by centralized training, which can result in more robust policies [59]. However, it is crucial to recognize that maintaining a centralized controller comes with its own set of challenges. Firstly, it can be prohibitively expensive and infeasible in certain scenarios, particularly in large-scale or resource-constrained environments. Additionally, the centralized controller introduces privacy concerns, as it necessitates the sharing of sensitive information among agents. Moreover, the central controller becomes a single point of failure, rendering the entire system vulnerable to disruptions. These trade-offs highlight the need for careful consideration of the CTDE framework’s suitability in specific MARL applications, taking into account the advantages and drawbacks it presents.
Figure 4: Illustration of centralized training with decentralized execution (CTDE) in MARL.
II-C2 Decentralized training with decentralized execution (DTDE)
In decentralized settings, individual agents acquire knowledge independently, devoid of direct interactions. Each agent operates with its distinct set of observations, policies, and algorithms, utilizing the environment as the exclusive channel for engagement. Commonly employed algorithms within this framework encompass independent Q-learning (IQL) [54] and independent advantage actor-critic (IA2C) [4], frequently adopted as foundational benchmarks. Although these algorithms enable autonomous learning by agents, they may not fully harness the collaborative potential [60]. To address non-stationarity concerns, [61] proposes a fully decentralized MARL framework. In this framework, each agent makes autonomous decisions based on both its locally observed environment and the messages exchanged with neighboring agents via the network, offering convergence guarantees. Fig. 5 shows the DTDE pattern based on the actor-critic framework.
Figure 5: Illustration of decentralized training with decentralized execution (DTDE) in MARL.
II-D MARL algorithm variants
In this subsection, we have exclusively reviewed the common MARL algorithms that find utility in CAV applications. Further, the MARL algorithms are divided into four categories: value function decomposition, learning to communicate, hierarchical structure, and causal inference, which are rooted in the inherent complexity and diverse requirements of multi-agent environments, particularly in CAV control. Table I offers an extensive overview of key works from four distinct perspectives within various settings.
TABLE I: An overview of the primary works for MARL algorithms.
Categories Work MARL algorithm                 Novelty
Value Function Decomposition [62], 2018 VDN Decomposing value functions using linear summation
[63], 2018 QMIX Enforcing monotonicity constraints between joint and individual value functions
[64], 2019 QTRAN Relaxing the non-negative weight network in Qmix
[65], 2020 WQMIX Introducing weighting functions and non-monotonic mixing networks
[66], 2021 AVD-Net An attention-based approach to capitalize on the coordination relations between agents
[67], 2023 PER Both the reward of an agent acting on its own and cooperating with other observable agents
Learning to Communicate [68], 2017 FingerPrint Merging low-dimensional policy fingerprints in the state space of neighboring agents
[69], 2016 DIAL The agent continuously encodes information and sends it to other recipients
[33], 2020 NeurComm Encoding and concatenating communicating signals rather than aggregating them
[70], 2020 IMAC Learning an efficient communication protocol based on the information bottleneck principle
Hierarchical Structure [71], 2019 FHM Establishing a feudal hierarchy structure
[72], 2018 HQMIX Including hierarchical QMIX and hierarchical communication network
[73], 2019 HSD Training a cooperative decentralized policy for high-level skill selection
[74], 2019 RODE Introducing a method to clearly divide the action space through action clustering
[75], 2023 HAVEN Designing a dual coordination mechanism within a two-level hierarchical structure
Causal Inference [76], 2018 CIR Agents receive a reward based on their causal influence on the actions of others
[77], 2023 ICL Evaluating the causal effect of each agent’s observations on team collaboration performance
[78], 2023 LAIES Calculating the causal effect of their actions on external states using the do-calculus process
[79], 2022 FD-MARL Constructing continuous communication protocols based on causal analysis
[80], 2022 DVD Deconfounded value function decomposition based on causal effects
II-D1 Value function decomposition
The problem of credit assignment in a fully cooperative environment has recently become a research hotspot. The shared rewards in a fully cooperative environment make it difficult to distinguish the contribution of each agent, some agents tend to be lazy [81]. To solve the above problem, some studies learn different value functions for each agent to distinguish the contribution of each agent. Foerster et al. [82] and Guo et al. [83] introduced the counterfactual baseline principle to learn different value functions by centralized learning method in the context of a shared team reward environment. Hou et al. utilized a role attention credit allocation mechanism to manage the role policy learning process in multi-agent systems [84]. Liu et al. construct the causal effect of the agent’s actions on the external state from the perspective of a causal diagram, solving the problem of lazy agents [78].
Among the above methods to solve credit assignment, value function decomposition is considered to be an important research direction. The main idea of value function decomposition is to decompose a joint value function into individual value functions to distinguish the different contributions of each agent. In the process, many useful value function methods were proposed to solve increasingly complex problems. VDN decomposed the joint action-value function into the sum of individual action-value functions for each agent, this linear decomposition method can achieve good results on some simple problems [62]. Different from the linear value function decomposition method of VDN, QMIX imposes monotonicity constraints between the joint value function and the individual value function, which enhances the representation ability of the joint value function and can deal with more and more complex scenarios [63]. QTRAN further improves the representation ability of joint functions by relaxing the non-negative weight network in QMIX [64]. WQMIX extends the value function method to complex scenarios of non-monotonic rewards by introducing weighting functions and non-monotonic mixing networks [65]. After that, some variant algorithms based on VDN and QMIX were used to solve different specific problems [66, 67].
II-D2 Learning to Communicate
MARL research can be categorized into two groups based on their communication methods [33]. The first group operates without communication among agents and centers its efforts on stabilizing training by employing advanced value estimation methods. For example, MADDPG [85] extends DDPG to the MARL setting by employing a centralized critic network to enable the global value estimate. Similarly, COMA [86] extends the actor-critic algorithm to the MARL setting by estimating each advantage using a centralized critic and a counterfactual baseline. The second group explores heuristic communication protocols, which may involve either direct message sharing or learnable communication protocols. For instance, in FingerPrint [68], the low-dimensional policy fingerprints are directly shared and incorporated into the state space of neighboring agents. In DIAL [69], a communicated message is generated simultaneously with the estimation of action values by each DQN agent. This message is then encoded and aggregated with other input signals at the receiver side. NeurComm proposed in [33] argues that encoding and concatenating communicating signals, rather than aggregating them, offers distinct advantages in mitigating information loss during communication. IMAC method learns an efficient communication protocol as well as scheduling based on the information bottleneck principle [70]. Their method involves the development of a novel differentiable communication protocol, which incorporates information related to state, policy, and the encoded data from neighboring agents. Subsequently, this information is encoded and concatenated into the state space of the neighboring agents.
II-D3 Hierarchical Structure
In the realm of multi-agent systems, a plethora of challenges are encountered, encompassing issues like sparse rewards, sequential decision-making, and limited transfer capabilities. To address these challenges effectively, the adoption of a hierarchical approach is unequivocally advantageous. The fundamental tenet of hierarchical multi-agent reinforcement learning revolves around the acquisition of hierarchical strategies for decision-making across varying levels of temporal abstraction. Current research in this domain broadly categorizes hierarchical multi-agent reinforcement learning into two paradigms: option-based hierarchical structures [87, 88] and goal-based hierarchical structures [89, 90], both of which have found applications in diverse multi-agent scenarios.
Feudal Multi-agent Hierarchies (FHM) establish a feudal hierarchy within a multi-agent context, yielding noteworthy results. However, a primary limitation of this approach is its inherent challenge in tackling fully collaborative tasks with shared rewards [71]. To address the predicament of multi-agent scenarios characterized by sparse and delayed rewards, several hierarchical structures with temporal abstraction have been proposed, including hierarchical QMIX and hierarchical communication network methods [72].
Furthermore, recognizing that certain hierarchical structures tend to heavily rely on manually crafted high-level action spaces, hierarchical learning with skill discovery (HSD) is designed to train a cooperative decentralized policy for high-level skill selection, coupled with the training of independent low-level policies for skill execution. This approach enhances adaptability to dynamic scenarios [73]. RODE, on the other hand, introduces a method for clear action space partitioning through action clustering, where each action corresponds to a distinct subspace, this innovation substantially mitigates the challenge of extensive parameter tuning associated in HSD [74]. Moreover, HAVEN introduces a dual coordination mechanism, encompassing inter-layer and inter-agent strategies, facilitated by the formulation of a reward function within a two-level hierarchical structure. This approach effectively addresses the instability issue through the concurrent optimization of strategies at all levels and inter-agent coordination [75]. Hierarchical multi-agent reinforcement learning is recognized as a promising avenue for addressing collaborative decision-making challenges in complex, large-scale scenarios. Nonetheless, it faces the challenge of difficult to design complex dynamic hierarchical structures and difficult migration of policies.
II-D4 Causal Inference
There are multiple complex variables in multi-agent reinforcement learning, and it is usually difficult to directly discover their internal relationships. Therefore, some recent research has begun to combine MARL with causal inference to discover the causal relationships between agents or variables, and further understand the operating mechanism of agents through intervention and inference, thereby motivating the agent to conduct more targeted learning [91]. Pearl’s three-level causal model is considered a powerful tool for constructing causal relationships between variables [92]. Causal influence reward methods promote the collaborative performance of agents by rewarding those agents that have a causal influence on the actions of other agents, where this causal influence is evaluated by counterfactual reasoning [76]. Rafael Pina et al proposed the independent causal learning (ICL) algorithm to evaluate the causal effect of each agent’s observations on team collaboration performance and solve the credit allocation problem in the independent learning framework [77].
There are also other works that combine causal inference from different perspectives under the framework of MARL. LAIES mathematically define the concept of fully lazy agents and teams by calculating the causal effect of their actions on external states using the do-calculus process, which solved the sparse reward problem in MARL [78]. FD-MARL allows agents to modify communication messages by choosing the counterfactual that bears the most significant influence on others, the continuous communication based on causal analysis enables efficient information transformation in a fully decentralized manner [79]. The deconfounded value decomposition (DVD) method investigates value function decomposition from the perspective of causal inference, which cuts off the backdoor confounding path from the global state to the joint value function [80]. In addition, causal inference has also been applied to some practical application scenarios, such as traffic signal control [93] and human-computer interaction decision-making [94].
TABLE II: An overview of the primary works from 1-D and 2-D cooperation for MARL applied in CAV control.
Scenarios Typical application Work MARL algorithm Novelty    Metrics
1-D Cooperation Platooning control [95], 2020 LSTM-Based MA-Reinforce $\bullet$ Individual and global rewards design $\bullet$ A LSTM-based communication protocol $\bullet$ Convergence stability $\bullet$ Travel time
[96], 2020 MA-DDQN $\bullet$ Local channel data incorporation $\bullet$ A different reward design including the weighted sum-rates and the transmission time $\bullet$ Sum-rate of V2N $\bullet$ Packet probability of V2V
[33], 2020 NeurComm $\bullet$ A spatial discount factor to stabilize training $\bullet$ A new differentiable communication protocol on both agent states and behaviors $\bullet$ Average headway $\bullet$ Average velocity
[97], 2021 C-HARL $\bullet$ Cooperative hierarchical attention RL (C-HARL) $\bullet$ Meta-learning with GNN integration $\bullet$ Convergence $\bullet$ Success rate of message $\bullet$ Communication delay
[98], 2022 MADDPG $\bullet$ A Stackelberg game to model interactions $\bullet$ Decisions sharing with social partners within the platoon due to social effects $\bullet$ Convergence $\bullet$ Defined utility value $\bullet$ Social effect
[99], 2023 Distributed PPO $\bullet$ State fusion strategy with the equilibrium to stabilize traffic oscillations $\bullet$ A reward function design in quadratic form $\bullet$ Driving comfort $\bullet$ Stability $\bullet$ Travel efficiency
2-D Cooperation Cooperative lane changing [100], 2021 MAPPO $\bullet$ A decentralized cooperative lane-changing design $\bullet$ A novel reward function design considering eliminating traffic shock waves $\bullet$ Traffic throughput $\bullet$ Number of Stops $\bullet$ Fuel efficiency
[101], 2021 Zero-sum DQN $\bullet$ The delay of an individual vehicle and overall traffic efficiency in the reward function $\bullet$ A lane change model with limited sensing without V2X communications $\bullet$ Convergence $\bullet$ Stability $\bullet$ Travel efficiency
[102], 2021 MADQN $\bullet$ Interactions and negotiations model establishment between multiple agents $\bullet$ An experience replay mechanism to tackle the non-stationarity $\bullet$ Timeout $\bullet$ Accident $\bullet$ Success rate
[103], 2022 QMIX $\bullet$ The simplification of system complexity through reinforcement learning $\bullet$ The fairness in cooperation to allow for independent lane-changing and overtaking $\bullet$ Collaborative effect $\bullet$ Travel efficiency
[104], 2022 Bi-level DQN $\bullet$ An innovative reward function designed with their own benefits and the impact on overall traffic $\bullet$ The driving intentions of surrounding vehicles encoded into the observation space $\bullet$ Comfort $\bullet$ Safety $\bullet$ Travel efficiency
[105], 2022 MA2C $\bullet$ A novel local reward design and parameter-sharing $\bullet$ Multi-objective reward function design $\bullet$ Driving comfort $\bullet$ Safety $\bullet$ Travel efficiency
III Towards applications of MARL in CAVs
CAVs have garnered substantial interest for their potential to reshape the transportation landscape, ushering in an era marked by heightened efficiency, enhanced safety, and bolstered sustainability [106]. Operating within networked environments, CAVs engage in interactions with multiple vehicles, promoting cooperative driving. For example, CAVs can acquire the skills to collaborate, execute seamless merges, and navigate intersections with precision, thereby advancing traffic safety and optimizing flow [106]. These transformative capabilities are increasingly being explored in tandem with MARL to further optimize CAV interactions within networked environments. Within the realm of MARL, CAVs are designed to engage in intelligent interactions with multiple vehicles, thus promoting cooperative driving strategies.
In this section, we will comprehensively explore the recent strides made in the utilization of MARL within CAV applications. Our examination will be structured according to various dimensions of cooperation, each correlated with a specific number of control components.
  • One-dimensional cooperation corresponds to scenarios involving control along a single control direction, such as either longitudinal control or lateral control.
  • Two-dimensional cooperation extends the scope to include both longitudinal and lateral control components, reflecting the increased complexity in the coordination and decision-making of CAVs.
  • Three-dimensional cooperation further augments the challenges and opportunities by incorporating additional constraints, such as time limits, which encompass aspects like traffic light control and on-ramp merging.
By categorizing these advancements based on control components, we aim to provide a comprehensive perspective on the evolving landscape of MARL within CAV applications, with a focus on the intricacies of cooperation and control in varying dimensions. Additionally, we provide an extensive examination of the leading simulation platforms employed in MARL research for CAVs.
III-A One-dimensional cooperation
In this subsection, we will conduct a detailed review of one-dimensional cooperation applications, which primarily focus on longitudinal control. One prominent example of such application is platooning control, where CAVs operate in closely-knit formations and collaborate to enhance efficiency, safety, and fuel economy. This is achieved by minimizing aerodynamic drag and optimizing traffic flow through synchronized control strategies and inter-vehicle coordination. Platooning control stands as a noteworthy illustration of how MARL contributes to the advancement of transportation systems.
As shown in Fig. 6, the platooning control problem is often formulated as a model-free multi-agent network [33], where each agent (i.e., AV) is capable of communicating with the vehicles ahead and behind via V2V communication channels, to achieve safe, fuel efficiency, and smooth vehicle following while maximizing the benefits of close formation driving. They learn cooperative strategies to adapt their behaviors based on the actions of other agents in the network.
Figure 6: One-dimensional cooperation scenarios: the case of platooning control.
One critical task of platooning control is to achieve string stability, which refers to the ability of a line of CAVs to maintain a stable and orderly formation (e.g., stable distance and speed) as they travel in close proximity to each other [107]. Notably, MARL has been proposed to address and enhance the achievement of string stability in the platooning control scenario, contributing to the development of more efficient and coordinated platooning control strategies. For instance, in [95], a MARL approach based on a robust communication protocol with Long Short-Term Memory (LSTM) [108] is introduced, resulting in stable platooning control. In addition, Li et al. introduce the Communication Proximal Policy Optimization (CommPPO) algorithm to enhance stability, which adapts to varying agent numbers and supports different dynamics of platooning control. CommPPO incorporates a predecessor-leader-follower communication protocol to facilitate the transmission of both global and local state information among agents. Notably, CommPPO introduces a novel reward communication channel, effectively mitigating issues related to spurious rewards and mitigating the problem of “lazy agents”, which are commonly encountered in other MARL approaches [109]. In [33], the platooning control problem is formulated as a spatiotemporal Markov decision process. They enhance system stability by introducing a spatial discount factor for local agents. Furthermore, they introduce a novel differentiable communication protocol, NeurComm, which mitigates non-stationarity, leading to improved learning efficiency and control performance.
Efficient communication among CAVs in platooning control also poses a fundamental challenge in achieving seamless coordination and maintaining the desired following properties, such as precise inter-vehicle spacing and synchronized maneuvers. For instance, Liu et al. introduce a Multi-Agent Hierarchical Attention Reinforcement Learning (MAHARL) framework, which incorporates the Graph Attention Network (GAT) at each hierarchical level to account for the mutual influences among agents. The hierarchical architecture of MAHARL empowers agents with foresight, enabling them to make favorable decisions, even in the absence of immediate incentives, and consider long-term rewards, as elaborated in their work [97]. Additionally, in scenarios where complete information is lacking, Li et al. [98] devise a game-theoretic approach to model the strategic interactions between the leading vehicle and the following vehicles. To overcome the absence of full information, the equilibrium solution in platooning control is identified through backward induction, coupled with information collection from the following vehicles.
Mixed platooning control refers to the coexistence of multiple platoons. In this context, it is crucial for each platoon to possess comprehensive knowledge about the other existing platoons and vehicles in the network, including both inter-platooning connectivity, which pertains to communication between different platoons, and intra-platoon connectivity, which relates to communication between vehicles within a platoon [110]. Parvini et al. present two innovative MARL algorithms, namely, modified MADDPG and modified MADDPG with task decomposition. In these algorithms, each platoon leader acts as an independent agent, engaging with the environment to learn its optimal policy. Notably, these algorithms employ multiple critics to estimate both global and local rewards, thereby fostering cooperative behavior among the agents. In the second algorithm, individual rewards are further decomposed into task-specific sub-reward functions, offering insights into their work as detailed in [111].
Even though CAVs have made significant advancements, there will be an extended period during which CAVs will coexist with human-driven vehicles (HDVs), a situation commonly referred to as mixed-traffic scenarios [4]. Different penetration rates can be assessed to understand the impact of mixed traffic. In the study presented in [112], the impact of the penetration rate of CAVs on the energy efficiency of the traffic network is examined, with a specific emphasis on a cooperative eco-driving system that utilizes longitudinal control. A noteworthy aspect of this research is the introduction of a role transition protocol for CAVs, enabling them to smoothly transition between leading and following positions within a vehicle string. Lu et al. enhance the evaluation by integrating several key elements, including altruism control, a quantitative car-following strategy, a refined platoon reward function, and a collision avoidance method, as outlined in their work [113]. In [99], the authors introduce the concept of characterizing consecutive HDVs as a collective entity, referred to as AHDV, to minimize stochastic variability and leverage macroscopic characteristics for the control of following CAVs. They employ a control strategy built on distributed proximal policy optimization (DPPO) to anticipate disturbances and downstream traffic conditions in mixed traffic scenarios. In vehicular environments characterized by high mobility, the use of traditional centralized optimization methods that rely on global channel information can be impractical. In [96], the authors model each CAV as an individual agent, which makes decisions based on local information and communication with neighboring vehicles, without relying on a centralized controller.
III-B Two-dimensional cooperation
Compared to one-dimensional cooperation, two-dimensional cooperation addresses problems related to both longitudinal and lateral control, with its most typical application being cooperative lane changing. As shown in Fig. 7, CAVs communicate and coordinate through V2V and V2I communication channels, enabling them to make decisions and acquire effective lane-changing strategies within shared driving environments.
Figure 7: Two-dimensional cooperation scenario.
Hou et al. devise a decentralized cooperative lane-changing controller employing a Multi-Agent Proximal Policy Optimization (MAPPO) approach, which empowers each vehicle to independently learn and assess its policy and action rewards based on local information, while still having access to global state data. The trained policies can be effectively transferred and applied across various traffic conditions, enhancing traffic throughput from uncongested to highly congested scenarios [100]. In [103], the authors put forth a twin-vehicle cooperative driving approach based on the QMIX algorithm. Their method harnesses RL for real-time adaptation in dynamic highway scenarios, striking a balance between autonomy and cooperation via a fair cooperation algorithm, which enables independent lane changing and overtaking in heavy traffic conditions and ensures a fixed formation. These capabilities enable the twin vehicles to boost their speed while skillfully avoiding rigid traffic interference. An enhanced QMIX algorithm [114] is further proposed to enhance the flexibility and effectiveness of the collaborative lane-changing system. They incorporate a stable estimation technique to mitigate the influence of overestimated joint Q-values among agents, which finely balance between maintaining formation and permitting unhindered overtaking, thereby enabling intelligent adaptation to diverse scenarios including heavy traffic, loose traffic, and emergency situations. In an effort to enhance mobility within complex traffic environments, in [115], instead of relying on relative distance and semantic maps, the authors suggest the utilization of traffic states that encapsulate the spatio-temporal interactions between neighboring vehicles. Within the MADRL framework, three prediction models, namely the transformer-based (TS-Transformer), generative adversarial network-based (TS-GAN), and conditional variational autoencoder-based (TS-CVAE) models, are developed and compared for traffic state prediction.
Graph neural networks (GNNs) [116] have gained significant attention in MARL settings, particularly for cooperative lane-changing applications. In the study by Chen et al. [117], they introduce a novel algorithm based on MADQN, which combines a GCN with a deep Q-network. This approach facilitates effective information fusion and decision-making within the MARL framework, thereby enhancing both safety and mobility in cooperative lane-changing scenarios. Moreover, in [118], the authors leverage GNN in conjunction with MADDPG in their work for multi-agent training, addressing the complexity of real-world inputs and aiding in congestion mitigation efforts.
Safety is of paramount importance in the context of CAV scenarios. In [119], the authors have introduced a safety-enhancing actor-critic algorithm that incorporates two innovative techniques. The first technique is the ‘truncated Q-function’, which effectively leverages shared information from neighboring CAVs, ensuring scalability for large-scale CAV systems. The second technique, “safe action mapping”, provides safety guarantees throughout both the training and execution phases by utilizing control barrier functions. Additionally, a bi-level strategy is proposed in [104] to further enhance safety and efficiency. At the upper level, the MADQN model is employed to make lane-change decisions, taking into account the collaborative nature of driving and considering the driving intentions of surrounding vehicles, enabling implicit negotiation for right-of-way. At the lower level, a right-of-way assignment model is utilized to ensure safety. A novel reward function is further proposed to encourage coordination and account for traffic impact. Moreover, Li et al. have introduced MetaDrive, a platform that focuses on safe driving by addressing generalizability, safety awareness, and multi-agent decision-making. It offers diverse driving scenarios for benchmarking single and multi-agent RL tasks [120].
Thoughtful reward function design is paramount for the effectiveness and efficiency of lane-changing strategies. In [105], the authors introduce the multi-agent advantage actor-critic (MA2C) framework, which integrates a novel local reward design and parameter-sharing scheme for cooperative lane changing in mixed traffic scenarios. They also present a multi-objective reward function that considers factors such as fuel efficiency, driving comfort, and safety to facilitate successful and efficient lane changing. In [101], a MARL approach is explored in which agents collaborate to reach a zero-sum game state. In contrast to cooperative driving, this approach, known as harmonious driving, places emphasis on achieving a balance between overall and individual efficiency while utilizing limited sensing data from individual vehicles. They design a reward function that promotes harmony by considering both individual and overall efficiency, as opposed to competitive strategies that solely prioritize individual interests.
III-C Three-dimensional cooperation
Three-dimensional cooperation involves coordinating both longitudinal and lateral motion as well as the temporal dimension in diverse traffic scenarios, including traffic signal coordination, on-ramp merging, and unsignalized intersections, as depicted in Fig. 8. To ensure safe, efficient, and effective cooperation, CAVs need to rapidly learn longitudinal and lateral maneuvers within tight time constraints.
Figure 8: Three-dimensional cooperation scenarios: (a) Traffic Signal Coordination; (b) Merging on-ramps;(c) Unsignalized intersections.
III-C1 Traffic signal control (TSC)
TSC aims to alleviate congestion in saturated road networks by adapting signal timings based on real-time traffic conditions (see Fig. 8 (a)). Traditional TSC methods address congestion by solving optimization problems to determine effective coordination and control policies. However, this process can be time-consuming and often requires precise system dynamics [4]. Recently, model-free MARL approaches have shown a growing interest in precise TSC.
Value-based MARL algorithms have been extensively investigated for TSC challenges. For instance, to alleviate the curse of dimensionality and environmental nonstationarity problems, a decentralized coordination MADQN approach is proposed in [121] to explicitly identify and dynamically adapt agent coordination needs during the learning process. Similarly, in [122], the complex TSC problem is decomposed into simpler subproblems and tackled using multiple regional agents and a centralized global agent. Each regional agent learns its RL policy for smaller regions with reduced action spaces, while the centralized global agent combines the RL contributions from various regional agents to form a final Q-function for the entire large-scale traffic grid. Notably, QMIX, an extension of VDN, was investigated in [123] for TSC, achieving promising results. Additionally, a multi-agent dueling-double-deep Q network (MAD3QN) is introduced in [124] for TSC. This framework incorporates a novel $\gamma$-reward design that includes both the original $\gamma$-reward and $\gamma$-attention-reward. By employing a spatial differentiation method for agent coordination, their approach enables decentralized control and decoupling of road networks, enhancing scalability and convergence. In [125], a cooperative group-based multi-agent Q-learning for ATSC (CGB-MATSC) is proposed. To enhance the learning ability, k-nearest-neighbor-based state representation, pheromone-based regional green-wave control mode, and spatial discounted reward design are introduced.
On the other hand, policy-based MARL algorithms have also been widely studied for TSC. In [32], a fully scalable and decentralized Multi-Agent Actor-Critic (MA2C) algorithm is introduced, which incorporates an advanced discount factor to mitigate the effects of remote agents. This approach demonstrates significant advantages over random and independent controllers. Wu et al. introduced a novel multi-agent recurrent deep deterministic policy gradient (MARDDPG) algorithm for TSC in vehicular networks [126]. They adopt a strategy of centralized training and decentralized execution, while also implementing parameter sharing among the actor networks to accelerate the training process. Additionally, LSTM networks are incorporated to leverage historical information for improved control. In [127], a MA2C approach with a feudal hierarchy concept is presented. In this approach, the traffic network is divided into several regions, each monitored by a manager agent along with traffic signal agents as workers. Managers handle high-level coordination and goal-setting for workers who control traffic signals to achieve these goals. This hierarchical structure allows for global coordination while preserving scalability.
Efficient communication is crucial in applying MARL to TSC as it enables coordinated decision-making, optimal resource allocation, adaptation to dynamic environments, scalability, conflict resolution, and efficient learning, all of which contribute to effective traffic management [128]. Liu et al. present a novel algorithm that enhances communication efficiency through a new message exchange method and improves congestion measurement by introducing a more comprehensive reward calculation method. Then a clear and simple representation of traffic conditions is provided, and the synchronization between agents at different intersections with varying cycle lengths needs to be addressed [129].
Graph-based MARL offers advantages in flexible representation, scalability, efficient message passing, a global view, interpretable representations, and adaptability, making it suitable for various multi-agent scenarios, including TSC problems [130]. In [131], the authors present CoLight, a model that employs graph-attentional MARL networks to enable communication among traffic signals, integrating temporal and spatial influences from neighboring intersections to a target intersection while establishing index-free modeling of these neighboring intersections. In [132], an Inductive Heterogeneous Graph Multi-agent Actor-Critic (IHG-MA) algorithm for multi-intersection traffic signal control is proposed, featuring two key aspects. Firstly, it utilizes an inductive heterogeneous graph (IHG) neural network for representation learning, effectively encoding both heterogeneous features and structural information of nodes and graphs to generate embeddings for previously unseen nodes and new traffic networks. Secondly, IHG-MA adopts the MA2C framework for policy learning, utilizing these embeddings to compute Q-values and policies.
To account for the spatial influence of multi-intersection traffic lights and the temporal dependency on historical traffic conditions, Wang et al. introduce a novel spatio-temporal MARL framework (STMARL) [133]. This framework captures spatio-temporal dependencies among multiple traffic lights, coordinating their control through a traffic light adjacency graph constructed based on spatial structure and integrating historical traffic records with current traffic status using a recurrent neural network. Additionally, a graph neural network model is designed to represent relationships among multiple traffic lights based on temporally-dependent traffic information. Considering the importance of shared information, relying on heuristic methods. Jiang et al. propose a new communication form called UniComm, which incorporates numerous observations gathered at one intersection and predicts their impact on neighboring intersections, enhancing communication efficiency [134]. Antonio et al. leverage the multi-agent TD3 algorithm by incorporating an LSTM to address the issue of varying observation shapes based on the number of vehicles, and then employ a training method called curriculum through self-play to enable collaborative control of CAVs at intersections [135].
III-C2 On-ramps merging
Highway on-ramp merging is a challenging task for AVs, as it involves the efficient merging of on-ramp CAVs onto the through lane without collisions (see Fig. 8 (b)). In this process, CAVs on the through lane need to proactively adjust their speed, either decelerating or accelerating, to create sufficient space for on-ramp CAVs to merge safely. Simultaneously, on-ramp CAVs adapt their speed and execute lane changes promptly when it is safe to do so, in order to prevent deadlock situations [136, 35].
In [137], a simplified mathematical formulation for an on-ramp merging scenario is presented to model the fundamental interactions and solved with a MADQN algorithm that accounts for the interaction between an on-ramp merging vehicle and a traffic vehicle in the target lane. The results indicate that a multi-agent approach can result in reduced collision rates compared to a single-agent approach, but this improvement is contingent on the optimality of the traffic vehicle’s behavior in the target lane. Zhou et al. introduce a cooperative merging control strategy for CAVs using a distributed MADDPG approach, which accounts for safe merging distances, acceleration limits, and factors like rear-end safety, lateral safety, and energy consumption [138]. To address the challenge of a dynamic environment resulting from the decentralized learning of CAVs, a decentralized framework employing MADDPG is presented to coordinate CAVs during highway merging [139]. This framework enables policies learned from a small group of trained CAVs to be transferred and applied to any number of CAVs, with a reward function that promotes high-speed travel, resulting in smoother traffic flow while maintaining safety in terms of rear-end and lateral collisions.
Some studies also examine the influence of human drivers in on-ramp merging scenarios. For instance, Hu et al. introduce interaction-aware decision-making with an adaptive strategies (IDAS) approach, which utilizes a modified MA2C method and curriculum learning to train a single policy that enables an AV to navigate the road while taking into account the cooperativeness of other drivers [140]. In [35], the authors present MA2C approach that enables CAVs on both the merge lane and through the lane to collaboratively learn a policy, maximizing traffic throughput while adapting to the behavior of HDVs. Their efficient and scalable MARL framework accommodates dynamic traffic scenarios with varying communication topologies, utilizing parameter sharing, local rewards, and action masking to promote inter-agent cooperation and scalability. Additionally, a priority-based safety supervisor is introduced to reduce collision rates and expedite the training process. Similarly, a game-theoretic multi-agent planning approach called GAMEPLAN is introduced in [141], which incorporates both human drivers and CAVs in merging scenarios. This method leverages game theory and an auction mechanism to compute optimal actions for each agent, considering their driving style inferred from commonly available sensors. The algorithm assigns higher priority to more aggressive or impatient drivers and lower priority to more conservative or patient drivers, ensuring game-theoretic optimality while preventing collisions and deadlocks.
TABLE III: An overview of the primary works from four perspectives in different settings.
Scenarios Typical application Work MARL algorithm Novelty    Metrics
3-D Cooperation Traffic signal control [121], 2019 MADQN $\bullet$ The state space of CAV decomposed into independent and coordinated parts $\bullet$ Dynamic adaptation to mitigate the curse of dimensionality and nonstationarity $\bullet$ Travel efficiency $\bullet$ Intersection delay $\bullet$ Decision quality
[124], 2021 MADQN $\bullet$ The $\gamma$-reward method to extend the Markov chain to the space-time domain $\bullet$ The spatial differentiation method to amend the current reward by recursion $\bullet$ Asymmetry road network evaluation $\bullet$ Attention score $\bullet$ Training time
[125], 2021 CG-Based MA- Q Learning (CGB-MAQL) $\bullet$ Agents organized into cooperative groups with each agent responsible for its region $\bullet$ A k-nearest-neighbor-based joint state representation among neighboring agents $\bullet$ A heuristic training mechanism by terminating poor behavior strategies $\bullet$ Waiting time $\bullet$ Road pheromone
[129], 2023 MA-Q Learning $\bullet$ A new message sending and processing $\bullet$ A data structure to record the latest and most valuable message $\bullet$ A new reward calculation with both queue length and total waiting time $\bullet$ Waiting time $\bullet$ Reward
On-ramps merging [137], 2019 MA-Q Learning $\bullet$ Multi-agent simulator proposed to study the interactions between a single pair of vehicles $\bullet$ A simplified mathematical formulation to capture the complexities of merging scenarios $\bullet$ Collision rates $\bullet$ Average speed
[139], 2022 MADDPG $\bullet$ Transferring the policies through a few learning CAVs in the training process $\bullet$ Reward function design at high speeds $\bullet$ Traffic flow smoothness
[138], 2022 Distributed MADDPG $\bullet$ The interference of human-driven Vehicles (HDV) $\bullet$ Reward function design including rear-end safety, lateral safety, and energy consumption $\bullet$ Fuel consumption $\bullet$ Travel time
[35], 2023 MA2C $\bullet$ An action masking scheme $\bullet$ Parameter sharing and local rewards to encourage inter-agent cooperation $\bullet$ A novel priority-based safety supervisor introduced to reduce collision rates $\bullet$ Traffic throughput $\bullet$ Collision rate
Unsignalized intersections [142], 2021 MA-Q Learning $\bullet$ The central section (CS) and the waiting section (WS) to introduce a regulation scheme $\bullet$ Features incorporation such as fixed Q-targets and experienced replay $\bullet$ Traffic throughput $\bullet$ Waiting time
[143], 2022 MADDQN $\bullet$ Route-agents introduced to coordinate the actions of multiple vehicles $\bullet$ The inclusion of a collision term in the reward function enables cooperation $\bullet$ The reuse of knowledge from agents’ policies to handle unknown traffic scenarios $\bullet$ Safety $\bullet$ Travel time $\bullet$ Fuel consumption
[144], 2022 QMIX $\bullet$ End-to-end learning to autonomously learn complex real-life traffic dynamics $\bullet$ Curriculum through Self-Play to adapt and learn in new and complex traffic scenarios $\bullet$ Travel time $\bullet$ Time lost due to congestion $\bullet$ Waiting time
[145], 2022 QMIX $\bullet$ Network-level improvements and Q value updates using Temporal Difference (TD) ($\lambda$) $\bullet$ Reward clipping operations to enhance the QMIX $\bullet$ Average speed $\bullet$ Collision rate $\bullet$ Fuelconsumption
III-C3 Unsignalized intersections
Intersections serve as pivotal nodes and potential bottlenecks within urban road networks. Enhancing traffic efficiency at these junctions is essential in ameliorating overall traffic throughput and alleviating congestion. Nevertheless, the management of vehicular interactions at unsignalized intersections, as illustrated in Fig. 8 (b), presents a notably intricate challenge [143]. Many conventional techniques, such as rule-based [146], planning-based [147], and single-agent RL methods [148], tend to approach the problem of intersection management as if it were a single-agent challenge, which can not capture the complex interactions among CAVs [149]. Recently, MARL has emerged as a promising and effective tool for managing traffic at unsignalized intersections.
In [143], an advanced collaborative multi-agent double Deep Q-Network (DQN) framework is proposed for unsigned traffic management. Their framework incorporates an efficient reward function that accounts for both the safety and efficiency of CAVs. Furthermore, this framework offers adaptability, enabling transfer learning and the reuse of knowledge from agent policies, particularly when dealing with unfamiliar traffic scenarios. In [150], the authors introduce a multi-agent delayed-A3C (MAD-A3C) approach, which utilizes continuous, model-free DRL to train a neural network capable of predicting acceleration and steering angles of other road users at each time step. Their study illustrates that these CAV agents can acquire the essential rules for effectively navigating intersections by comprehending the priorities of other learners in the environment, all the while ensuring safe driving along their designated routes.
Yan et al. utilize multi-agent policy decomposition, enabling decentralized control based on local observations for any number of controlled CAVs. Remarkably, even without reward shaping, the DQN-based approach learns to coordinate CAVs to exhibit traffic signal-like behaviors, achieving near-optimal throughput with 33%-50% controlled vehicles. Furthermore, through multi-task learning and transfer learning, they demonstrate the generalizability of this behavior across different inflow rates and traffic network sizes [151].
Effective coordination among CAVs is essential to ensure safe and efficient maneuvers at unsignalized intersections. In [152], a decentralized MADQN algorithm is adopted and learned to make decisions for CAVs. To enable effective coordination among agents, the intent trajectories of other neighboring agents are incorporated into each agent’s state space. Furthermore, a decentralized and conflict-free coordination scheme designed for CAVs at unsignalized intersections is proposed in [149], with the objective of enhancing intersection management precision. They frame the challenge of safely guiding multiple vehicles through unsignalized intersections as a partially observable stochastic game (POSG). To facilitate collaborative decision-making in a distributed fashion, they introduce a cooperative multi-agent proximal optimization algorithm (CMAPPO). In [153], a multi-layer coordination strategy is proposed for the management of unsignalized intersections. This architecture comprises two layers: a low-level layer that employs a dynamics-based algorithm to control individual CAVs, and a high-level layer that integrates a twin delayed deep deterministic policy gradient (TD3) algorithm, trained in a centralized manner but executed in a decentralized fashion by multiple agents for decision-making. Their findings reveal a successful training process, with the MATD3 algorithm achieving a remarkable 100% success rate in preventing intersection collisions. To tackle coordination between adjacent intersections, a multi-agent-based deep reinforcement learning scheduling (MA-DRLS) algorithm is proposed in [142]. This approach allows each intersection agent to independently devise an optimal scheduling strategy through information exchange with other agents. The algorithm employs DQN networks, and the use of fixed Q-targets and experience replay helps improve the neural network’s reliability during the training process. In [154], a multi-agent joint-action DDPG (MAJA-DDPG) framework that integrates learning and sequential optimization is proposed for the management of unsignalized intersections. The algorithm involves two primary steps: Initially, a shared policy is learned to determine the order in which vehicles cross at the intersection based on traffic state information. Subsequently, vehicle trajectories are optimized sequentially following the determined crossing order. Notably, the proposed algorithm learns a shared policy that can be implemented in a distributed manner.
To address the challenge of limited communication resources, the authors in [155] introduce an algorithm known as Efficient Communication Method (ECM)-MA2C. This approach aims to maintain coordination performance while making efficient use of constrained communication resources. The ECM-MA2C algorithm employs a variational auto-encoder algorithm combined with advanced multi-head attention mechanisms to extract and retain valuable information from neighboring agents. Experimental results demonstrate that this efficient communication scheme outperforms baseline approaches in bandwidth-constrained environments, highlighting its effectiveness in achieving improved performance with limited communication resources.
In [145], the authors have concentrated their efforts on the application of a value decomposition-based MARL approach, namely QMIX, for the coordination of multiple CAVs at unsignalized intersections. To enhance the performance of the original QMIX framework, several implementation enhancements have been integrated. These improvements encompass network-level optimizations, Q-value updates using Temporal Difference (TD) with $\gamma$, and reward clipping operations.
Figure 9: Representative scenarios in different simulators: (a) SUMO [156]; (b) CityFlow [157]; (c) SMARTS [34]; (d) MetaDrive [158]; (e) CARLA [159]; (f) MACAD [160].
III-D Simulation platforms
The simulation platforms, which act as the training/learning environments, play an important role in the performance and adaptability of the autonomous agents. The learning platforms provide different challenges for the planning and control of the autonomous agents by considering different traffic scenarios and different characters of agents (Fig. 9). These platforms offer a virtual environment to learn and validate the CAV’s perception, decision-making, and low-level control systems, where the cost and the safety of the agents can be optimized. The traffic ecosystem/simulation platform incorporates realistic models, physics-based simulations, and efficient numerical solvers, providing a controllable environment to accelerate the development of CAVs [161]. Comparisons of different simulation platforms for CAVs are shown in Table IV.
SUMO (Simulation of Urban MObility) is an open-source traffic simulation package used to model CAVs and their interactions [156]. It provides tools for simulating traffic scenarios, including vehicle behavior and road networks, enabling researchers to study and analyze traffic dynamics. SUMO aids in evaluating traffic management strategies, routing algorithms, and the impact of CAVs on traffic flow and efficiency. It is widely used for assessing CAV behavior and optimizing transportation systems in a virtual environment.
CityFlow, by optimizing data structures and using efficient algorithms, improves the simulation efficiency and accelerates the learning tasks, allowing for large-scale road networks and traffic flow simulations [157]. It supports flexible definitions of road networks using synthetic and real-world data and offers a user-friendly interface for reinforcement learning. CityFlow outperforms SUMO, being over twenty times faster and capable of city-wide traffic simulations with interactive rendering. It opens new possibilities for testing machine learning methods in the intelligent transportation domain, serving as a foundation for transportation studies beyond traffic signal control.
To mitigate the gaps between learning and reality, learning platforms with more realistic agent characters are required, and several simulation platforms considering vehicle dynamics and realistic sensor data are proposed.
SMARTS (Scalable Multi-Agent RL Training School) is a dedicated simulation platform for realistic multi-agent interaction in autonomous driving research [34]. SMARTS supports the training and accumulation of diverse behavior models, enabling the creation of realistic and varied driving interactions. It offers a user-friendly interface, key features, and benchmark tasks. The platform is open-source, encouraging research on multi-agent learning for autonomous driving.
MetaDrive stands out as an efficient and compositional driving simulator, offering infinite scene generation, lightweight operation, realistic physics simulation, and multiple sensory inputs [158], where a trade-off between the visual rendering and the physical simulation has been made. With the ability to generate infinite scenes encompassing diverse road maps and traffic settings, researchers can explore a wide range of driving environments, paving the way for the development of generalizable RL algorithms. These features make it a valuable tool for researchers in autonomous driving and robotics, enabling them to conduct impactful studies with comprehensive realism and flexibility.
CARLA, which is an open-source simulator developed by the team at Intel Labs, considers more comprehensive environment and agent variances. It provides a realistic 3D environment for testing and evaluating autonomous driving algorithms, including support for multi-agent scenarios [159]. With CARLA, researchers and developers can create complex simulations that closely resemble real-world driving conditions, enabling them to assess and refine the performance of their autonomous driving systems. Its high-fidelity physics engine and accurate sensor models allow for realistic perception and control experiments. Furthermore, CARLA offers extensive customization options, facilitating the creation of various urban environments, weather conditions, and traffic scenarios to comprehensively evaluate the robustness and effectiveness of autonomous driving algorithms.
MACAD (Multi-Agent Connected and Autonomous Driving), as another high-fidelity simulation platform for CAVs, adopts "partially observable Markov games" to formulate the connected autonomous driving problems with realistic assumptions and modeling. The platform provides an extensible set of CAD simulation environments that enable the research and development of Deep RL-based integrated sensing, perception, planning, and control algorithms for CAD systems with unlimited operational design domains under realistic, multi-agent settings [160]. The MACAD also provides a taxonomy of multi-agent learning environments based on the nature of tasks, the nature of agents, and the nature of the environment, which can be tuned by the researchers.
More simulators for CAVs have been designed to consider more detailed characters and versatile scenarios, which provides more tools to mitigate the gaps between the simulation and reality.
A PCMA (pedestrian crash avoidance mitigation) system is developed for interdependent decisions of the vehicles and pedestrians by Trumpp et. al [162]. In the system, an AV-pedestrian interaction at an unmarked crosswalk is modeled using a Markov decision process and deep reinforcement learning (DRL). Two pedestrian behaviors are considered: a baseline predefined strategy and an advanced model using DRL, turning the interaction into a multi-agent problem. The PCAM systems are assessed based on collision rates and traffic flow efficiency, focusing on the impact of observation uncertainty.
Nocturne considers the limited visibility for studying the multi-agent coordination in the 2D driving simulator, as the complex partial observablilities of the multi-agent system also attract the attention of the researchers. It efficiently calculates visible features, allowing for rapid simulations. Using real-world driving data, tests showed that machine-learning agents significantly differed from human-like coordination [163].
Generally speaking, the simulation-based evaluations are simpler to implement, reproduce, and scale compared with real experiments for learning. However, simulations may not capture all the challenges associated with an actual deployment. For example, factors such as network delay, vehicle model discrepancies, computation time, and the necessity of implementing clock synchronization and fail-safe routines pose challenges in real-world implementations.
TABLE IV: Comparasions between some typical CAV simulation platforms.
Simulators
Vehicle
Dynamics
3D
Scenarios
Visual
Rendering
Pedestrians
Light
Weighted
Customized
Maps
Data
Importing
Hightway-env [164]
Nocturne [163]
SUMO [156]
CityFlow [157]
SMARTS [34]
MetaDrive [158]
CARLA [159]
MACAD [160]
IV Discussions & Future Works
CAVs leveraging MARL algorithms present a promising frontier in the evolution of transportation. As urban centers grapple with traffic congestion and road safety, the nuanced and adaptive behaviors offered by MARL algorithms can enable CAVs to respond more fluidly to the actions of other road participants. The dynamism of MARL allows these vehicles to cooperatively optimize traffic flows, potentially alleviating congestion and reducing transit times. Nevertheless, the journey to this future is not without its challenges. Interactions with unpredictable human drivers, managing the delicate balance between exploration and safety in learning, and navigating a complex regulatory landscape are just a few of the hurdles. Yet, with ongoing collaborations among tech giants, startups, and academic circles, there is a collective push to overcome these barriers, signaling an optimistic trajectory for the marriage of MARL and CAVs.
IV-A Macro-micro energy optimization of CAVs using MARL
Efficient energy management in autonomous vehicles is crucial, as it not only ensures the optimal performance of each vehicle but also contributes to overall energy conservation [165]. This can be divided into two distinct yet complementary approaches: microscopic energy optimization, which concentrates on maximizing the efficiency of individual components within a single vehicle, and macroscopic energy optimization, which aims to enhance collective energy use across a fleet of autonomous vehicles.
Microscopic energy optimization scrutinizes the performance of each component, aiming for optimal functioning. Traditional methods, which segregate powertrain, chassis control, and other controls, overlook the synergistic potential between these elements, potentially missing out on energy savings. To overcome this, Hua et al. present a collaborative strategy between two learning agents within the MADRL framework [166]. They conduct a sensitivity analysis to harmonize the settings for two learning agents, identifying an optimal mode that surpasses single-agent models by saving up to 4% more energy. Yang et al. develop a multi-objective energy management strategy for hybrid electric vehicles, leveraging MARL [167]. Their strategy simultaneously addresses fuel economy, battery charge maintenance, battery longevity, and ultracapacitor constraints, considering the engine–generator set (EGS) and hybrid energy storage system (HESS) as interactive agents. Shi et al. propose an energy management system (EMS) that uses independent Q-learning, a multi-agent reinforcement learning algorithm, to sustain battery charge levels and minimize hydrogen consumption, thereby optimizing energy for HESS[168]. Microscopic energy optimization enhances component efficiency but risks overlooking vehicle-wide dynamics. Macroscopic optimization addresses this by aiming to improve the energy efficiency across a fleet of autonomous vehicles.
Macroscopic energy optimization has garnered increasing interest, particularly concerning traffic dynamics and their impact on energy efficiency. For instance, Wang et al. have developed a MARL-based strategy for hybrid electric vehicles that synchronizes powertrain management with car-following behaviors, thereby reducing energy use while maintaining safe distances [169]. Zhang et al. have designed a MARL method that simultaneously optimizes velocity and energy management for groups of HEVs, employing Markov games and LSTM networks for dynamic group understanding and an innovative asynchronous learning technique for inter-vehicle knowledge sharing[170]. Furthermore, Peng et al. utilize a MADDPG framework with two distinct agents for adaptive cruise control and energy management, respectively, achieving domain-specific optimization and enhancing the MADDPG’s effectiveness by more than 10% with an improved prioritized experience replay technique [171].
Macro-micro energy optimization: The advancement of CAVs is set to enhance sustainable and efficient mobility, with macro-micro energy modeling playing a key role in precise energy consumption analysis, as depicted in Figure 10. This approach intertwines macro-level traffic analysis with micro-level hybrid electric vehicle (HEV) powertrain operations. By doing so, it allows for adaptive energy management that responds to real-time traffic conditions and powertrain status, optimizing the use of the electric motor and battery in various driving scenarios. Consequently, this integrated optimization strategy not only improves route efficiency and driving behavior but also leads to notable energy savings and cost reductions [172].
Figure 10: Illustration of the macro-micro energy management system control framework.
IV-B Communication challenges
Current progress: Communication in MARL for CAVs is a critical element that underpins the efficiency and privacy of these intelligent transportation systems. The ability of CAVs to collaborate, share information, and make coordinated decisions hinges on effective communication protocols [173]. First, the development of advanced communication protocols tailored specifically for CAVs is a pressing need. Existing research focuses on protocols that enable efficient data sharing [174]. For instance, research endeavors have delved into the utilization of graph-based MARL algorithms [117], such as those explored in [130] and [132], to enable efficient and adaptive message exchange among agents. Second, the challenge of ensuring robust communication in dynamically evolving traffic scenarios represents another critical aspect that deserves meticulous attention [175]. For example, in [155], a variational auto-encoder algorithm combined with advanced multi-head attention mechanisms is proposed to extract and retain valuable information from neighboring agents while making efficient use of constrained communication resources. Last but not least, it is worth noting that managing unstable communication channels, particularly in terms of latency, emerges as a particularly intriguing consideration in the context of MARL for CAVs [96]. In [96], the double deep Q-learning algorithm is proposed to jointly train the agents to maximize the sum rate of V2N links while ensuring the desired packet delivery probability for each V2V link within specified latency constraints.
Research gaps: While communication in MARL is a well-explored domain, its adaptation and application within CAVs present distinctive research gaps, which necessitate further exploration [128]. For instance, existing works often overlook pressing real-world concerns in the realm of CAVs, such as communication costs and the challenges posed by noisy environments. These critical factors, when ignored, can introduce substantial obstacles to the practical deployment of CAV technology. MARL algorithms applied in computer science domains are expected to be explored for use in CAV applications within challenging environments marked by limited bandwidth [176] and noisy communication channels [177]. Moreover, the majority of current MARL algorithms in CAVs rely on Centralized Training and Decentralized Execution (CTDE, Sec. II), potentially posing a significant privacy threat within the realm of communication. In the field of single-agent RL, security and privacy issues have garnered extensive attention [178]. However, in the context of MARL, and similarly in the domain of CAVs, these areas have remained relatively unexplored. This suggests a promising area for future research that merits exploration.
IV-C Mixed-traffic challenges
Current progress: One of the formidable challenges posed by CAVs stems from navigating the intricate dynamics of mixed traffic, wherein CAVs share the road with various other road users, including human-driven vehicles, bicycles, e-scooters, and an array of other modes of transportation [179, 180, 181]. This multifaceted coexistence necessitates advanced MARL algorithms to ensure seamless and safe interactions among these diverse road users. Current research into MARL algorithms for CAV applications primarily concentrates on integrating the prediction of human behavior into the decision-making process [140, 182]. For instance, in [35], the authors introduce a novel MARL framework with a priority-based safety supervisor designed to preempt potential collisions. This is achieved by predicting future trajectories, encompassing both CAVs and HDVs and then the unsafe actions generated by MARL agents will be replaced with safe actions, thereby contributing to the enhancement of safety in mixed traffic scenarios. However, it is important to note that human-driven vehicles can introduce an additional layer of unpredictability into the traffic environment, since mixed traffic involves a diverse range of road users, presenting unique challenges for the seamless integration of CAVs into the existing road ecosystem.
Research gaps: Given the diverse spectrum of road users, future research endeavors will encompass the development of real-time, adaptive safety assessment and intervention strategies capable of responding to the dynamic nature of traffic conditions. Another significant aspect entails the establishment of novel coordination and communication mechanisms between CAVs and other road users. For instance, as the presence of e-scooters on the roads continues to surge, there is a growing imperative to develop efficient methods for coordinating the interactions between CAVs and e-scooters [183]. Furthermore, the existing MARL algorithms are typically trained within specific scenarios, underscoring the necessity for research aimed at enhancing their capacity to generalize across a wide range of mixed traffic conditions and environmental contexts.
IV-D Sim-to-real challenges
Deep MARL has achieved significant breakthroughs in various areas of the CAV domain. Given the constraints associated with real-world data collection, such as sample inefficiency and cost considerations, simulation environments have become invaluable for agent training (see Sec. III-D). Nevertheless, the disparity between the simulated and real-world settings can lead to a degradation in policy performance when models are transitioned to actual CAV scenarios [36]. In certain simulator platforms, a key objective is to enable the deployment of CAV agents with minimal disparities between their training data and real-world experiences [32, 11]. Simultaneously, other research initiatives are dedicated to enhancing safety, recognizing it as a primary obstacle to achieving real-world online training for complex self-driving car agents [35]. Conversely, it is crucial to explore advanced MARL techniques, including approaches like multi-agent offline RL [184] that leverage pre-existing offline data to update policies with online adaptations. Such future exploration can open new avenues for enhancing the generalization abilities and performance of CAVs in real-world scenarios.
V Conclusion
In recent times, MARL has emerged as a focal point of interest within the realm of ITS. The capability of MARL to tackle intricate control and coordination challenges in CAVs has unlocked new horizons for the development of intelligent and efficient next-generation transportation systems. This review has undertaken a comprehensive exploration of the applications of MARL in the context of CAVs.
  1. 1.
    The review began with an overview of single-agent RL techniques and an extensive survey of the diverse landscape of MARL architectural variants.
  2. 2.
    Our analysis categorized these contributions based on different degrees of control, ranging from one-dimensional to two-dimensional and three-dimensional collaboration. Within this multifaceted arena, MARL has exhibited remarkable prowess, demonstrating its promising performance in various CAV control tasks. These include cooperative endeavors such as adaptive and cruise control, lane-changing maneuvers, on-ramp merging, traffic light coordination, and unscripted interactions, among others.
  3. 3.
    Amidst these accomplishments, we have also illuminated the formidable challenges that lie ahead. Designing and evaluating MARL methodologies within the CAV environment present their own set of complexities and intricacies. Balancing performance, scalability, and safety is no small feat.
This review stands as a valuable resource, poised to catalyze further research and the practical application of MARL within the domains of CAVs and ITS.